# Task-Augmented Cross-View Imputation Network for Partial Multi-View Incomplete Multi-Label Classification

### v1

@@ III-C Cross-View Imputation
After obtaining the embedding features of each view using VSEncoder, to fully utilize and effectively complement the unique information of each view, we adopt a straightforward and efficient strategy, namely multi-view weighted fusion [9][18], which systematically merges all views through weight assignment to form a unified and comprehensive representation as follows:
@@
Learning this comprehensive sample representation $\widehat{Z}$ by exploiting the shared nature of view-level information described above means that if some views are absent from the sample, a symmetric cross-view decoder module VSDecoder connected to $\widehat{Z}$ can be trained to reconstruct the data for all views, especially targeting the originally missing parts for completion, i.e.:
@@
III-D Incomplete Multi-Label Classification
Once the complete views $\left\{T_{l}\right\}_{l=1}^{m}$ have been obtained, they are then used for the final, incomplete multi-label classification task. Firstly, pre-trained view-specific encoders $\left\{\Psi_{l}\right\}_{l=1}^{m}$ are employed to capture the deep semantic information in each $T_{l}$:
@@ III-D Incomplete Multi-Label Classification
Then, our target is to create a unified and consistent representation which adequately characterizes multi-view data. However, different views may make varying contributions to such a representation. To account for the different discriminative capabilities of different views, and the relative values explorability of individual view under the data completeness enhancement, we transition to a weighted fusion approach with learnable weights as follow when the originally missing views are made complete to substitute the fixed-weight fusion method employed when multi-view data are incomplete [42][43]:
@@
Finally, we construct a direct mapping from the joint feature space to the multi-label classification space. It is achieved by applying two linear layers with a sigmoid activation function that independently generates a probability score for each potential category in each sample. These probability scores are bounded by a sigmoid transformation between 0 and 1. Mathematically, this classification process can be abstracted as $\left\{\Theta:\widehat{Z_{T}}\in R^{n\times d_{e}}\longrightarrow Y^{{}^{\prime}}\in\left[0,1\right]^{n\times c}\right\}$, where $Y^{{}^{\prime}}$ represents the probability distribution of $c$ categories predicted for $n$ samples, $\Theta$ denotes the classifier. And to address the issue of missing labels, we employ a missing label indicator matrix, denoted by $G$, to adjust the BCE loss function [18][30]. This result in the proposal of a weighted binary cross-entropy loss function, which assesses the difference between the predicted label $Y^{{}^{\prime}}$ and the real label $Y$. This assessment guides the model training by following:
$L_{c}$ $=\frac{1}{nc}\sum_{i=1}^{n}\sum_{j=1}^{c}\left[Y_{i,j}\log{Y_{i,j}^{{}^{\prime}}}\right.$ (13)
$\left.+\left(1-Y_{i,j}\right)\log{\left(1-Y_{i,j}^{{}^{\prime}}\right)}\right]G_{i,j}$
@@
  Input: Partial multi-view data $\left\{X_{l}\right\}_{l=1}^{m}$, accompanied by a missing-view indicator matrix $U\in\left\{0,1\right\}^{n\times m}$; Incomplete multi-label matrix $Y\in R^{n\times c}$, accompanied by a missing-label indication matrix $G\in\left\{0,1\right\}^{n\times c}$; batch size $B$; learning rate $l_{r}$; hyper-parameters $\alpha$ and $\left\{\delta_{l}\right\}_{l=1}^{m}$.
  Initialization: Fill in missing elements with 0, initialize TAEncoder {϶$\left\{\backepsilon_{l}\right\}_{l=1}^{m}$, VSEncoder $\left\{\Psi_{l}\right\}_{l=1}^{m}$, VSDecoder $\left\{\Phi_{l}\right\}_{l=1}^{m}$, classifiers $\left\{\complement_{l}\right\}_{l=1}^{m}$ and $\Theta$.
@@
     calculate $V_{l}=\backepsilon_{l}\left(X_{l}\right)$, $Y_{l}^{{}^{\prime}}=\complement_{l}\left(V_{l}\right)$;calculate $L_{ta}^{l}$ by Eq.(6);
     updata ϶$\backepsilon_{l}$ and $\complement_{l}$;
@@
  for $l=1$ to $m$ do
     calculate multi-view embedding feature $Z_{l}$ by Eq.(7);calculate fusion feature $\widehat{Z}$ by Eq.(8);calculate reconstructed feature $V_{l}^{{}^{\prime}}$ by Eq.(9);calculate reconstructed loss $L_{re}$ by Eq.(10);complete missing data in $X_{l}$ with $V_{l}^{{}^{\prime}}$ to get new data $T_{l}$;calculate multi-view embedding feature $Z_{T}^{l}$ by Eq.(11);calculate fusion feature $\widehat{Z_{T}}$ by Eq.(12);calculate classification loss $L_{c}$ by Eq.(13);
     updata $\Psi_{l}$, $\Phi_{l}$, $\Theta$ and $\sigma_{l}$;
@@ IV Experiment
IV-A Experimental Settings
IV-A1 Datasets
In accordance with previous studies [11, 17, 44], we have chosen five widely used multi-view multi-label datasets to verify our proposed model. These datasets include Corel5k, Pascalo7, ESPGame, IAPRTC12, and MIRFLICKR. Each dataset encompasses six views, specifically GIST, HSV, HUE, SIFT, RGB, and LAB. A brief summary of these datasets can be found in Table II, with further detailed descriptions provided as follows:
TABLE II: Statistics of five multi-view multi-label datasets.
@@ IV-A1 Datasets
IV-A2 Incomplete Multi-view Multi-label Data Preparation
@@
Given that all five datasets mentioned above have complete views and labels, they cannot be used directly to evaluate the efficacy of our model in situations with missing data. Following the methodology in [11, 17, 44], we preprocess the complete dataset into a partial multi-view incomplete multi-label dataset. More specifically, we establish samples with a randomly missing percentage, denoted as $a$%, in each view, ensuring that each sample is present in at least one view, thereby creating a partial multi-view dataset. We then randomly eliminate $b$of both positive and negative labels for each category to simulate a scenario with missing labels. Finally, the dataset is randomly divided into three subsets: 70% for training, 15% for validation, and 15% for testing. Note that any missing element in views and labels is substituted with 0.
IV-A3 Competitors
@@
In the experimental session, we compare our TACVI-Net with four classical models and five recent popular approaches: GLOCAL [50], CDMM [22], iMvML [11], NAIM3L [17], DICNet [44], AIMNet [51], MTD [31], SIP [33], and ATSINet [32]. Since GLOCAL and AIMNet have not been discussed in detail in Section II-B, a brief description of these two methods is necessary here. GLOCAL emphasizes the exploration of label correlations from both global and local views. AIMNet is an attention-induced missing instance filling technique based on an attempt to approximate potential features of missing instances in the embedding space through cross-view joint attention. To address the incomplete multi-label classification property of GLOCAL, which is only applicable to single views, we conduct experiments independently for each view and select the best result to characterize the multi-view classification performance. In addition, as CDMM lacks the capacity to address missing data, we make suitable corrections to the model by implementing a mean-filling strategy. Our experiment is conducted ten times repetitively, with both the mean and standard deviation of the results being documented.
IV-A4 Evaluation metrics
@@
Consistent with prior studies [11, 17, 44], we employ four commonly utilized evaluation metrics for multi-label classification to assess our model’s performance, specifically: Average Precision (AP), Hamming Loss (HL), Ranking Loss (RL), and Area Under the Receiver Operating Characteristic Curve (AUC). To render the metrics more intuitively illustrative of the model’s performance, we convert HL into its inverse (1-HL) and similarly invert RL to (1-RL), where a higher score for each metric denotes superior performance.
IV-A5 Implementation details
@@
Our TACVI-Net model is implemented utilizing Python programming language and the Pytorch framework. For model training, we employ the Stochastic Gradient Descent (SGD) optimizer, with a momentum value set to 0.9. The configuration includes a batch size of 128 and a learning rate of 0.0001. The computational environment for running our model comprises an Ubuntu operating system, equipped with an RTX3080Ti GPU and an i-2700KF CPU.
@@
TABLE III: The experimental results of ten methods on five datasets with 50% missing views, 50% missing labels, and 70% training samples. The higher the value, the better the performance. Bold indicates the best result.
@@
DATA METRIC GLOCAL CDMM iMvWL NAIM3L DICNet AIMNet MTD SIP ATSINet TACVI-Net
AP 0.285$\pm$0.004 0.354$\pm$0.004 0.283$\pm$0.007 0.309$\pm$0.004 0.381$\pm$0.004 0.400$\pm$0.010 0.415$\pm$0.008 0.418$\pm$0.009 0.436$\pm$0.007 0.447$\pm$0.011
@@
AP 0.256$\pm$0.002 0.305$\pm$0.004 0.235$\pm$0.004 0.261$\pm$0.001 0.323$\pm$0.001 0.329$\pm$0.005 0.332$\pm$0.003 0.331$\pm$0.006 0.361$\pm$0.004 0.368$\pm$0.005
1-HL 0.980$\pm$0.000 0.981$\pm$0.000 0.969$\pm$0.000 0.981$\pm$0.000 0.981$\pm$0.000 0.981$\pm$0.000 0.981$\pm$0.000 0.980$\pm$0.000 0.981$\pm$0.000 0.981$\pm$0.000
IAPRTC12 1-RL 0.825$\pm$0.002 0.862$\pm$0.002 0.833$\pm$0.003 0.848$\pm$0.001 0.873$\pm$0.001 0.883$\pm$0.003 0.875$\pm$0.002 0.885$\pm$0.003 0.898$\pm$0.003 0.901$\pm$0.002
AUC 0.830$\pm$0.001 0.864$\pm$0.002 0.833$\pm$0.003 0.848$\pm$0.001 0.873$\pm$0.001 0.885$\pm$0.003 0.876$\pm$0.001 0.886$\pm$0.002 0.899$\pm$0.002 0.901$\pm$0.002
AP 0.537$\pm$0.002 0.570$\pm$0.002 0.495$\pm$0.012 0.551$\pm$0.002 0.589$\pm$0.005 0.602$\pm$0.004 0.607$\pm$0.004 0.614$\pm$0.004 0.631$\pm$0.003 0.633$\pm$0.003
1-HL 0.874$\pm$0.001 0.886$\pm$0.001 0.840$\pm$0.003 0.882$\pm$0.001 0.888$\pm$0.002 0.890$\pm$0.001 0.891$\pm$0.001 0.891$\pm$0.001 0.895$\pm$0.001 0.895$\pm$0.001
MIRFLICKR 1-RL 0.832$\pm$0.001 0.856$\pm$0.001 0.806$\pm$0.011 0.844$\pm$0.001 0.863$\pm$0.004 0.873$\pm$0.002 0.875$\pm$0.002 0.877$\pm$0.002 0.887$\pm$0.001 0.888$\pm$0.001
AUC 0.828$\pm$0.001 0.846$\pm$0.001 0.794$\pm$0.015 0.837$\pm$0.001 0.872$\pm$0.001 0.849$\pm$0.004 0.861$\pm$0.001 0.862$\pm$0.002 0.860$\pm$0.003 0.873$\pm$0.002
IV-B Experimental Result and Analysis
In Table III, we present the experimental outcomes for ten different methods applied to five datasets, under conditions where 50% of instances are missing, 50% of labels are absent, and only 70% of the samples are utilized for training. And the standard deviation is indicated by the numeral in the lower right corner of each reported result. To investigate the effect of varying levels of missing data in terms of views and labels on our model’s performance, we maintain one missing rate at 50% while adjusting the other missing rate to a series comprising 30%, 50%, and 70%. The resultant outcomes are documented in Fig.2. In addition, we extend our experiments to datasets with view and label missing rates both of 30%, 50% and 70%, respectively, as shown in Fig.3. From the analysis of Table III, Fig.2, Fig.3 and Fig.4, several key observations can be readily deduced:
@@ IV-B Experimental Result and Analysis
    Based on the consolidated analysis presented in Table III, our proposed methodology significantly outperforms all current approaches on the majority of selected evaluation metrics, validating its effectiveness and marking a remarkable lead. Furthermore, techniques such as DICNet and ATSINet demonstrate superior performance compared to the first four methodologies, thanks to their carefully designed deep learning architectures that are explicitly tailored to tackle the intricacies of the PMVIMLC task. This observation underscores the exceptional potential of deep learning in unraveling the inherent complexities associated with this specific challenge, and highlights its ability to tackle such complicated problems.
@@
Although our TACVI-Net chiefly targets the challenge of PMVIMLC, we extend our evaluation to encompass the model’s performance within a full-view and full-label setting. The results of this assessment are presented in Table IV. Notably, the model retains its superior performance under MVMLC, indicating that our approach is not only efficacious for PMVIMLC but also holds applicability within MVMLC.
(a) missing-view ratios
(b) missing-label ratios
(a) 30% view-missing ratio and 70% label-missing ratio
(b) 70% view-missing ratio and 30% label-missing ratio
(a) Corel5k
(b) Pascal07
TABLE IV: The experimental results of different methods on five data sets with full views and labels. The higher the value, the better the performance. Bold indicates the best result.
DATA METRIC GLOCAL CDMM iMvWL NAIM3L DICNet MTD ATSINet TACVI-Net
AP 0.386$\pm$0.006 0.489$\pm$0.004 0.313$\pm$0.005 0.327$\pm$0.004 0.509$\pm$0.002 0.521$\pm$0.010 0.543$\pm$0.012 0.567$\pm$0.011

### v2

@@ 3.3 Cross-View Imputation
After obtaining the embedding features of each view using VSEncoder, to fully utilize and effectively complement the unique information of each view, we adopt a straightforward and efficient strategy, namely multi-view weighted fusion [37], which systematically merges all views through weight assignment to form a unified and comprehensive representation as follows:
@@
Learning this comprehensive sample representation $\widehat{Z}$ by exploiting the shared property of view-level information described above means that if some views are absent from the sample, a symmetric cross-view decoder module VSDecoder connected to $\widehat{Z}$ can be trained to reconstruct the data for all views, especially targeting the originally missing parts for completion, i.e.:
@@
3.4 Incomplete Multi-Label Classification
Once the complete views $\left\{T_{l}\right\}_{l=1}^{m}$ have been obtained, they are then used for the final, incomplete multi-label classification task. Firstly, view-specific encoders $\left\{\Psi_{l}\right\}_{l=1}^{m}$ are employed to capture the deep semantic information in each $T_{l}$:
@@ 3.4 Incomplete Multi-Label Classification
Then, our target is to create a unified and consistent representation which adequately characterizes multi-view data. However, different views may make varying contributions to such a representation. To account for the different discriminative capabilities of different views, and the relative values explorability of individual view under the data completeness enhancement, we transition to a weighted fusion approach with learnable weights as follow when the originally missing views are made complete to substitute the fixed-weight fusion method employed when multi-view data are incomplete [47]:
@@
Finally, we construct a direct mapping from the joint feature space to the multi-label classification space. It is achieved by applying two linear layers with a sigmoid activation function that independently generates a probability score for each potential category in each sample. These probability scores are bounded by a sigmoid transformation between 0 and 1. Mathematically, this classification process can be abstracted as $\left\{\mathbb{C}:\widehat{Z_{T}}\in R^{n\times d_{e}}\longrightarrow Y^{{}^{\prime}}\in\left[0,1\right]^{n\times c}\right\}$, where $Y^{{}^{\prime}}$ represents the probability distribution of $c$ categories predicted for $n$ samples, $\mathbb{C}$ denotes the classifier. And to address the issue of missing labels, we employ a missing label indicator matrix, denoted by $G$, to adjust the BCE loss function [36][48]. This result in the proposal of a weighted binary cross-entropy loss function, which assesses the difference between the predicted label $Y^{{}^{\prime}}$ and the real label $Y$. This assessment guides the model training by following:
$L_{c}=-\frac{1}{nc}\sum_{i=1}^{n}\sum_{j=1}^{c}\left[Y_{i,j}\log{Y_{i,j}^{{}^{\prime}}}+\left(1-Y_{i,j}\right)\log{\left(1-Y_{i,j}^{{}^{\prime}}\right)}\right]G_{i,j}$ (13)
@@
Input: Partial multi-view data $\left\{X_{l}\right\}_{l=1}^{m}$, accompanied by a missing-view indicator matrix $U\in\left\{0,1\right\}^{n\times m}$; Incomplete multi-label matrix $Y\in R^{n\times c}$, accompanied by a missing-label indication matrix $G\in\left\{0,1\right\}^{n\times c}$; batch size $B$; maximum number of iterations $T$; learning rate $l_{r}$; hyper-parameters $\alpha$ and $\left\{\delta_{l}\right\}_{l=1}^{m}$.
Initialization: Fill in missing elements with 0, initialize TAEncoder {϶$\left\{\backepsilon_{l}\right\}_{l=1}^{m}$, VSEncoder $\left\{\Psi_{l}\right\}_{l=1}^{m}$, VSDecoder $\left\{\Phi_{l}\right\}_{l=1}^{m}$, classifiers $\left\{\complement_{l}\right\}_{l=1}^{m}$ and $\mathbb{C}$.
@@
     calculate $V_{l}=\backepsilon_{l}\left(X_{l}\right)$, $Y_{l}^{{}^{\prime}}=\complement_{l}\left(V_{l}\right)$;
     calculate $L_{ta}^{l}$ by Eq.(6);
     update ϶$\backepsilon_{l}$ and $\complement_{l}$;
@@
for $t=1$ to $T$ do
     calculate multi-view embedding feature $Z_{l}$ by Eq.(7);
     calculate fusion feature $\widehat{Z}$ by Eq.(8);
     calculate reconstructed feature $V_{l}^{{}^{\prime}}$ by Eq.(9);
     calculate reconstructed loss $L_{re}$ by Eq.(10);
     complete missing data in $X_{l}$ with $V_{l}^{{}^{\prime}}$ to get new data $T_{l}$;
     calculate multi-view embedding feature $Z_{T}^{l}$ by Eq.(11);
     calculate fusion feature $\widehat{Z_{T}}$ by Eq.(12);
     calculate classification loss $L_{c}$ by Eq.(13);
     update $\Psi_{l}$, $\Phi_{l}$, $\mathbb{C}$ and $\sigma_{l}$;
@@ 4 Experiment
4.1 Experimental Settings
Table 2: Statistics of five multi-view multi-label datasets.
@@ 4.1 Experimental Settings
4.1.1 Datasets
In accordance with previous studies [12][16][23], we have chosen five widely used multi-view multi-label datasets to verify our proposed model. These datasets include Corel5k, Pascalo7, ESPGame, IAPRTC12, and MIRFLICKR. Each dataset encompasses six views, specifically GIST, HSV, HUE, SIFT, RGB, and LAB. A brief summary of these datasets can be found in Table 2, with further detailed descriptions provided as follows:
4.1.2 Incomplete Multi-view Multi-label Data Preparation
@@
Given that all five datasets mentioned above have complete views and labels, they cannot be used directly to evaluate the efficacy of our model in situations with missing data. Following the methodology in [12][16][23], we preprocess the complete dataset into a partial multi-view incomplete multi-label dataset. More specifically, we establish samples with a randomly missing percentage, denoted as $a$%, in each view, ensuring that each sample is present in at least one view, thereby creating a partial multi-view dataset. We then randomly eliminate $b$of both positive and negative labels for each category to simulate a scenario with missing labels. Finally, the dataset is randomly divided into three subsets: 70% for training, 15% for validation, and 15% for testing. Note that any missing element in views and labels is substituted with 0.
4.1.3 Competitors
@@
In the experimental session, we compare our TACVI-Net with four classical models and five recent popular approaches: GLOCAL [54], CDMM [28], iMvML [16], NAIM3L [23], DICNet [12], AIMNet [48], MTD [37], SIP [39], and ATSINet [38]. Since GLOCAL and AIMNet have not been discussed in detail in Section 2.2, a brief description of these two methods is necessary here. GLOCAL emphasizes the exploration of label correlations from both global and local views. AIMNet is an attention-induced missing instance filling technique based on an attempt to approximate potential features of missing instances in the embedding space through cross-view joint attention. To address the incomplete multi-label classification property of GLOCAL, which is only applicable to single views, we conduct experiments independently for each view and select the best result to characterize the multi-view classification performance. In addition, as CDMM lacks the capacity to address missing data, we make suitable corrections to the model by implementing a mean-filling strategy. Our experiment is conducted ten times repetitively, with both the mean and standard deviation of the results being documented. And the final experimental results are obtained by selecting the model parameters that performed best on the validation set during the optimization process and evaluating them on the test set.
4.1.4 Evaluation metrics
@@
Consistent with prior studies [12][16][23], we employ four commonly utilized evaluation metrics for multi-label classification to assess our model’s performance, specifically: Average Precision (AP), Hamming Loss (HL), Ranking Loss (RL), and Area Under the Receiver Operating Characteristic Curve (AUC). To render the metrics more intuitively illustrative of the model’s performance, we convert HL into its inverse (1-HL) and similarly invert RL to (1-RL), where a higher score for each metric denotes superior performance.
4.1.5 Implementation details
@@
Our TACVI-Net model is implemented utilizing Python programming language and the Pytorch framework. For model training, we employ the Stochastic Gradient Descent (SGD) optimizer, with a momentum value set to 0.9. The configuration includes a batch size of 128 and a learning rate of 0.0001. The computational environment for running our model comprises an Ubuntu operating system, equipped with an RTX3080Ti GPU.
@@
Table 3: The experimental results of ten methods on five datasets with 50% missing views, 50% missing labels, and 70% training samples. The higher the value, the better the performance. Bold indicates both the best and the suboptimal results.
@@
DATA METRIC GLOCAL [54] CDMM [28] iMvWL [16] NAIM3L [23] DICNet [12] AIMNet [48] MTD [37] SIP [39] ATSINet [38] TACVI-Net
AP 0.285$\pm$0.004 0.354$\pm$0.004 0.283$\pm$0.008 0.309$\pm$0.004 0.381$\pm$0.004 0.400$\pm$0.010 0.415$\pm$0.008 0.418$\pm$0.009 0.436$\pm$0.007 0.447$\pm$0.011
@@
AP 0.256$\pm$0.002 0.305$\pm$0.004 0.237$\pm$0.003 0.261$\pm$0.001 0.323$\pm$0.001 0.329$\pm$0.005 0.332$\pm$0.003 0.331$\pm$0.006 0.361$\pm$0.004 0.368$\pm$0.005
1-HL 0.980$\pm$0.000 0.981$\pm$0.000 0.969$\pm$0.000 0.980$\pm$0.000 0.981$\pm$0.000 0.981$\pm$0.000 0.981$\pm$0.000 0.980$\pm$0.000 0.981$\pm$0.000 0.981$\pm$0.000
IAPRTC12 1-RL 0.825$\pm$0.002 0.862$\pm$0.002 0.833$\pm$0.002 0.848$\pm$0.001 0.873$\pm$0.001 0.883$\pm$0.003 0.875$\pm$0.002 0.885$\pm$0.003 0.898$\pm$0.003 0.901$\pm$0.002
AUC 0.830$\pm$0.001 0.864$\pm$0.002 0.835$\pm$0.001 0.850$\pm$0.001 0.874$\pm$0.000 0.885$\pm$0.003 0.876$\pm$0.001 0.886$\pm$0.002 0.899$\pm$0.002 0.901$\pm$0.002
AP 0.537$\pm$0.002 0.570$\pm$0.002 0.490$\pm$0.012 0.551$\pm$0.002 0.589$\pm$0.005 0.602$\pm$0.004 0.607$\pm$0.004 0.614$\pm$0.004 0.631$\pm$0.003 0.633$\pm$0.003
1-HL 0.874$\pm$0.001 0.886$\pm$0.001 0.839$\pm$0.002 0.882$\pm$0.001 0.888$\pm$0.002 0.890$\pm$0.001 0.891$\pm$0.001 0.891$\pm$0.001 0.895$\pm$0.001 0.895$\pm$0.001
MIRFLICKR 1-RL 0.832$\pm$0.001 0.856$\pm$0.001 0.803$\pm$0.008 0.844$\pm$0.001 0.863$\pm$0.004 0.873$\pm$0.002 0.875$\pm$0.002 0.877$\pm$0.002 0.887$\pm$0.001 0.888$\pm$0.001
AUC 0.828$\pm$0.001 0.846$\pm$0.001 0.787$\pm$0.012 0.837$\pm$0.001 0.849$\pm$0.004 0.861$\pm$0.001 0.862$\pm$0.002 0.860$\pm$0.003 0.872$\pm$0.003 0.873$\pm$0.002
4.2 Experimental Result and Analysis
In Table 3, we present the experimental outcomes for ten different methods applied to five datasets, under conditions where 50% of instances are missing, 50% of labels are absent, and only 70% of the samples are utilized for training. And the standard deviation is indicated by the numeral in the lower right corner of each reported result. To investigate the effect of varying levels of missing data in terms of views and labels on our model’s performance, we maintain one missing rate at 50% while adjusting the other missing rate to a series comprising 30%, 50%, and 70%. The resultant outcomes are documented in Fig.2. In addition, we extend our experiments to datasets with view and label missing rates both of 30%, 50% and 70%, respectively, as shown in Fig.3. From the analysis of Table 3, Fig.2, Fig.3 and Fig.4, several key observations can be readily deduced:
@@ 4.2 Experimental Result and Analysis
    Based on the consolidated analysis presented in Table 3, our proposed methodology significantly outperforms all current approaches on the majority of selected evaluation metrics, validating its effectiveness and marking a remarkable lead. Furthermore, techniques such as DICNet and ATSINet demonstrate superior performance compared to the first four methodologies, thanks to their carefully designed deep learning architectures that are explicitly tailored to tackle the intricacies of the PMVIMLC task. This observation underscores the exceptional potential of deep learning in unraveling the inherent complexities associated with this specific challenge, and highlights its ability to tackle such complicated problems.
@@
Although our TACVI-Net chiefly targets the challenge of PMVIMLC, we extend our evaluation to encompass the model’s performance within a full-view and full-label setting. The results of this assessment are presented in Table 4. Notably, the model retains its superior performance under MVMLC, indicating that our approach is not only efficacious for PMVIMLC but also holds applicability within MVMLC.
Table 4: The experimental results of different methods on five data sets with full views and labels. The higher the value, the better the performance. Bold indicates the best result.
DATA METRIC GLOCAL [54] CDMM [28] iMvWL [16] NAIM3L [23] DICNet [12] MTD [37] ATSINet [38] TACVI-Net
AP 0.386$\pm$0.006 0.489$\pm$0.004 0.313$\pm$0.005 0.327$\pm$0.004 0.509$\pm$0.002 0.523$\pm$0.003 0.543$\pm$0.012 0.567$\pm$0.011